# FedSSA: Semantic Similarity-based Aggregation for Efficient Model-Heterogeneous Personalized Federated Learning

### v1

@@ 6.3 Case Studies
We test the robustness of FedSSA to Non-IIDness and client participation rates under model-heterogeneous FL settings.
6.3.1 Robustness to Non-IIDness
We study the robustness of FedSSA and FedProto to non-IIDness under setting $(N=100,C=10\%)$ on CIFAR-10 and CIFAR-100. We vary the number of classes seen by each client as $\{2,4,6,8,10\}$ on CIFAR-10 and $\{10,30,50,70,90,100\}$ on CIFAR-100. Figure 6 shows that FedSSA consistently outperforms FedProto, demonstrating its robustness to non-IIDness. As the non-IIDness decreases (the number of classes seen by each client increases), accuracy degrades as more IID local data enhances generalization and reduces personalization.
6.3.2 Robustness to Client Participant Rates
We also test the robustness of FedSSA and FedProto to client participant rates $C$ under $(N=100,C=10\%)$ on CIFAR-10 (non-IID: 2/10) and CIFAR-100 (non-IID: 10/100). We vary the client participant rates as $C=\{0.1,0.3,0.5,0.7,0.9,1\}$. Figure 7 shows that FedSSA consistently outperforms FedProto, especially on the more complicated CIFAR-100 dataset, demonstrating its robustness to changes in client participant rates. Besides, as the client participant rates increase, model accuracy drops as more participating clients provide more IID local data, which in turn, improves generalization and reduces personalization.
Figure 6: Robustness to Non-IIDness.
@@ 6.3.2 Robustness to Client Participant Rates
6.4 Ablation Studies
FedSSA consists of two core modules: 1) the semantic similarity-based header parameter aggregation module (denoted as ‘M1’) for local-to-global knowledge transfer, and 2) the adaptive parameter stabilization-based header parameter fusion module (denoted as ‘M2’) for global-to-local knowledge transfer. ‘M1’ is responsible for collaborative training across clients with heterogeneous local models. Thus, we conduct ablation experiments on ‘M2’. We discuss three cases: A) similar to LG-FedAvg, each client replaces its historical local header with the latest global header; B) each client only replaces the seen-class parameters of its historical local header with those of the latest global header; and C) each client follows FedSSA. We evaluate these three cases under model-heterogeneous settings.
Table 3 shows that Case-C achieves the best performance, followed by Case-A and Case-B, demonstrating the effectiveness of the proposed header parameter fusion with the adaptive parameter stabilization strategy. The reasons are that Case-C integrates the local and global knowledge of seen classes, but Case-B ignores seen-class local knowledge. Case-A not only omits seen-class local knowledge, but also includes unseen-class global knowledge that may disturb the classification of seen classes.
Table 3: Ablation study results.
N=10, C=100% N=50, C=20% N=100, C=10%
Case M1 M2 CIFAR10 CIFAR100 CIFAR10 CIFAR100 CIFAR10 CIFAR100
A 96.36 73.26 92.68 58.01 88.51 49.95
B 93.01 71.54 82.42 52.14 80.41 49.95
C 96.54 73.39 95.83 64.2 92.92 57.29
@@ 7 Conclusions and Future Work
In this paper, we propose a novel personalized heterogeneous federated learning framework named FedSSA to enhance the performance and efficiency of model-heterogeneous personalized federated learning. It consists of two core modules: 1) local-to-global knowledge transfer by semantic similarity-based header parameter aggregation and 2) global-to-local knowledge transfer by adaptive parameter stabilization-based header parameter fusion, both of them enhance the personalization of each client’s heterogeneous local model and stabilize classification boundaries. Theoretical analysis shows that FedSSA could converge over wall-to-wall time. Extensive experiments demonstrate that FedSSA achieves the best classification accuracy while incurring the lowest communication and computational costs. It is also robust to non-IIDness and changes in client participation rates.
In subsequent research, we will improve FedSSA from two aspects: 1) designing a more effective aggregation rule for aggregating header parameters by class, and 2) devising a smoother decay function for adaptive parameter stabilization.
@@ Appendix A Pseudo Code of FedSSA
             /* Semantic Similarity-based Aggregation */ $\theta^{t,s}=\frac{1}{|\mathcal{K}_{s}^{t}|}\sum_{k\in\mathcal{K}_{s}^{t}}\theta_{k}^{t,s}$;
@@
                   /* Adaptive Parameter Stabilization Strategy */ $\hat{\theta}_{k}^{t,s}=\theta^{t-1,s}+\mu^{t}\cdot\theta_{k}^{t-1,s},s\in S_{k}$; $\mu^{t}=\begin{cases}\mu_{0}\cdot\cos(\frac{t}{2T_{stable}}\pi),&t\leq T_{stable};\\ 0,&t>T_{stable};\end{cases}$
@@
Appendix B Detailed Settings and Results in Experiments
Layer Name CNN-1 CNN-2 CNN-3 CNN-4 CNN-5
Maxpool1 2$\times$2 2$\times$2 2$\times$2 2$\times$2 2$\times$2
Maxpool2 2$\times$2 2$\times$2 2$\times$2 2$\times$2 2$\times$2
FC1 2000 2000 1000 800 500
FC2 500 500 500 500 500
FC3 10/100 10/100 10/100 10/100 10/100
model size 10.00 MB 6.92 MB 5.04 MB 3.81 MB 2.55 MB
Figure 8: Mean test accuracy vs. communication rounds in model-heterogeneous scenarios.
Appendix C Proof for Lemma 2
@@
Appendix D Proof for Theorem 1
@@
Appendix E Proof for Theorem 2

### v3

@@ 6.3 Case Studies
6.3.1 Robustness to Non-IIDness (Class)
We study the robustness of FedSSA and FedProto to non-IIDness controlled by the number of seen classes one client holds under (N=100, C=10%) on CIFAR-10 and CIFAR-100. We vary the number of classes seen by each client as $\{2,4,\ldots,10\}$ on CIFAR-10 and $\{10,30,\ldots,100\}$ on CIFAR-100. Figure 7 shows that FedSSA consistently outperforms FedProto, demonstrating its robustness to non-IIDness. As the non-IIDness decreases (the number of classes seen by each client rises), accuracy degrades as more IID local data enhances generalization and reduces personalization.
6.3.2 Robustness to Non-IIDness (Dirichlet)
We also test the robustness of FedSSA and FedProto to more complex non-IIDness controlled by a Dirichlet($\gamma$) function under setting (N=100, C=10%) on CIFAR-10 and CIFAR-100. We vary $\gamma=\{0.1,\ldots,0.5\}$ on both datasets. Figure 8 presents that FedSSA also outperforms FedProto across all non-IIDnesses.
Figure 7: Robustness to Non-IIDness (Class).
Figure 8: Robustness to Non-IIDness (Dirichlet).
@@ 6.3.2 Robustness to Non-IIDness (Dirichlet)
6.3.3 Robustness to Client Participation Rates
We also test the robustness of FedSSA and FedProto to client participation rates $C$ under (N=100, C=10%) on CIFAR-10 (non-IID: 2/10) and CIFAR-100 (non-IID: 10/100). We vary the client participation rates as $C=\{0.1,0.3,\ldots,1\}$. Figure 9 shows that FedSSA consistently outperforms FedProto, especially on the more complicated CIFAR-100 dataset, demonstrating its robustness to changes in client participation rates.
@@ 7 Conclusions and Future Work
In this paper, we propose a novel personalized heterogeneous federated learning framework named FedSSA to enhance the performance and efficiency of model-heterogeneous personalized federated learning. It consists of two core modules: 1) local-to-global knowledge transfer by semantic similarity-based header parameter aggregation and 2) global-to-local knowledge transfer by adaptive parameter stabilization-based header parameter fusion, both of them enhance the personalization of each client’s heterogeneous local model and stabilize classification boundaries. Theoretical analysis shows that FedSSA could converge over wall-to-wall time. Extensive experiments demonstrate that FedSSA achieves the best classification accuracy while incurring the lowest communication and computational costs.
In subsequent research, we will improve FedSSA from two aspects: 1) designing a more effective aggregation rule for aggregating header parameters by class, and 2) extend FedSSA to broader real-world applications.
Acknowledgments
This research is supported in part by the National Science Foundation of China under Grant 62272252 & 62272253, the Key Research and Development Program of Guangdong under Grant 2021B0101310002, and the Fundamental Research Funds for the Central Universities; National Key R&D Program of China 2021YFF0900800 & NSFC No.92367202; the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP-2020-019); the RIE 2020 Advanced Manufacturing and Engineering (AME) Programmatic Fund (No. A20G8b0102), Singapore; the China Scholarship Council.
@@ Appendix A Pseudo Code of FedSSA
             /* Semantic Similarity-based Aggregation */
$\theta^{t,s}=\frac{1}{|\mathcal{K}_{s}^{t}|}\sum_{k\in\mathcal{K}_{s}^{t}}\theta_{k}^{t,s}$;
@@
                   /* Adaptive Parameter Stabilization Strategy */
$\hat{\theta}_{k}^{t,s}=\theta^{t-1,s}+\mu^{t}\cdot\theta_{k}^{t-1,s},s\in S_{k}$;
$\mu^{t}=\begin{cases}\mu_{0}\cdot\cos(\frac{t}{2T_{stable}}\pi),&t\leq T_{stable};\\ 0,&t>T_{stable};\end{cases}$
@@
Appendix B Proof for Lemma 2
@@
Appendix C Proof for Theorem 1
@@
Appendix D Proof for Theorem 2
@@
Appendix E Detailed Settings and Results in Experiments
Layer Name CNN-1 CNN-2 CNN-3 CNN-4 CNN-5
Maxpool1 2$\times$2 2$\times$2 2$\times$2 2$\times$2 2$\times$2
Maxpool2 2$\times$2 2$\times$2 2$\times$2 2$\times$2 2$\times$2
FC1 2000 2000 1000 800 500
FC2 500 500 500 500 500
FC3 10/100 10/100 10/100 10/100 10/100
model size 10.00 MB 6.92 MB 5.04 MB 3.81 MB 2.55 MB
Figure 10: Mean test accuracy after smooth processing vs. communication rounds in model-heterogeneous scenarios.
Appendix F Ablation Studies
FedSSA consists of two core modules: 1) the semantic similarity-based header parameter aggregation module (denoted as ‘M1’) for local-to-global knowledge transfer, and 2) the adaptive parameter stabilization-based header parameter fusion module (denoted as ‘M2’) for global-to-local knowledge transfer. We conduct ablation experiments for ‘M1’ and ‘M2’, respectively.
Table 4: Ablation study results for ‘M1’.
N=10, C=100% N=50, C=20% N=100, C=10%
Case M1 M2 CIFAR10 CIFAR100 CIFAR10 CIFAR100 CIFAR10 CIFAR100
A 96.30 72.20 94.83 60.95 91.27 45.83
B 96.42 72.69 95.10 62.53 91.85 50.71
C 96.54 73.39 95.83 64.20 92.92 57.29
Ablation for ‘M1’. We consider three cases for ‘M1’: A) LG-FedAvg aggregates whole local headers to generate the global header; B) LG-FedAvg further apply adaptive parameter stabilization (‘M2’) on clients; C) FedSSA with both ‘M1’ and ‘M2’. Table 4 presents that the adaptive parameter stabilization strategy further improves LG-FedAvg, but it still performs worse than FedSSA, indicating the necessity of combining ‘M1’ and ‘M2’.
Table 5: Ablation study results for ‘M2’.
N=10, C=100% N=50, C=20% N=100, C=10%
Case M1 M2 CIFAR10 CIFAR100 CIFAR10 CIFAR100 CIFAR10 CIFAR100
A 96.36 73.26 92.68 58.01 88.51 49.95
B 93.01 71.54 82.42 52.14 80.41 49.95
C 96.54 73.39 95.83 64.20 92.92 57.29
Ablation for ‘M2’. We use ‘M1’ and discuss three cases for ‘M2’: A) similar to LG-FedAvg, each client replaces its historical local header with the latest global header; B) each client only replaces the seen-class parameters of its historical local header with those of the latest global header; and C) each client follows FedSSA. We evaluate these three cases under model-heterogeneous settings.
Table 5 shows that Case-C achieves the best performance, followed by Case-A and Case-B, demonstrating the effectiveness of the proposed header parameter fusion with the adaptive parameter stabilization strategy. The reason is that Case-C integrates the local and global knowledge of seen classes, but Case-B ignores seen-class local knowledge. Case-A not only omits seen-class local knowledge but also includes unseen-class global knowledge that may disturb the classification of seen classes. Since Case-A exchanges the whole local or global header between the server and clients, it improves the generalization to all classes but compromises the personalization to local seen classes. Besides, Figure 11 shows that FedSSA with ‘M2’ (Case-C) indeed can mitigate parameter shaking and speed convergence compared with not using ‘M2’ (Case-B), especially on CIFAR-10.
Figure 11: Test accuracy varies as communication round for FedSSA without or with adaptive parameter stabilization strategy.